at a temp TK the pressure of 4g aegon in a bulb is p. the bulb is put in a bath having temp 50K higher than the first one.0.8g of argon has to be removed to maintain initial pressure. what is the value of T?
 
options 510,200,73,100  in kelvin

Grade:11

1 Answers

Bijaya Subhadarshini
5 years ago
 
P1 = n1.R.T1/V1

P1 = n2.R.T1 + 50/V1 

n1.T1 = n2.T1 + 50

T1. n1/n2 - 1= 50 

n1/n2 = m1/m2 

T1.m1/m2 - 1= 50 

T1.4.0 g / 3. 2 g - 1= 50 

T1.1.25 - 1.00= 50 

T1 = 50/ 0.25= 200 K 

T = T1 = 200 K
  • Was this answer helpful
  •  
  •  
  • 11